OpenStudy (a.somera32):

how do i find the value of the expression -> sin 1.7 csc 1.7?

TheSmartOne (thesmartone):

Hint: csc(x) = 1/sin(x)

OpenStudy (a.somera32):

im still confused? @TheSmartOne

TheSmartOne (thesmartone):

change csc1.7 in to sin and simplify it once you use the hint I gave above, you should see it more clearly :)

OpenStudy (a.somera32):

is x = 1.7?

OpenStudy (pythagoras123):

\[\sin1.7cosec1.7=\sin1.7 \times \frac{ 1 }{ \sin1.7 }=\frac{ \sin1.7 }{ \sin1.7 }=1\]

OpenStudy (a.somera32):

thank you!

TheSmartOne (thesmartone):

Do you understand? :P

OpenStudy (a.somera32):

yes :)

TheSmartOne (thesmartone):

ok, great :)